Hi, I have a problem and wouldn't even know where to start with it!! It is:
the stock price of a company has risen at the rate of $5 per month over the last year. On 1 January it was $75. Write an equation that shows this relationship.
Thanks,
Christa.

Answer by venugopalramana(3286)

Hi, I have a problem and wouldn't even know where to start with it!! It is:
the stock price of a company has risen at the rate of $5 per month over the last year. On 1 January it was $75.
LET US TRY TO GO STEP BY STEP AND SEE THE ANSWER
PRICE ON 1 JAN..........75
PRICE ON 1 FEB..THAT IS AFTER 1 MONTH.........75+5=75+1*5 = 80
PRICE ON 1 MAR..THAT IS AFTER 2 MONTHS........75+10=75+2*5 = 85
PRICE ON 1 APR..THAT IS AFTER 3 MONTHS........75+15=75+3*5 = 90
PRICE ON 1 MAY..THAT IS AFTER 4 MONTHS........75+20=75+4*5 = 95
....................................................................
.......................................................SIMILARLY...
PRICE AFTER N MONTHS .........................75+5N
..................................
PRICE ON 1 JAN NEXT YEAR ...AFTER 12 MONTHS...75+12*5=135
HENCE THE GENERAL EQUATION FOR PRICE AFTER N MONTHS =75+5N
Write an equation that shows this relationship.
Thanks,
Christa.